# Wi-Fi Radar via Over-the-Air Referencing: Bridging Wi-Fi Sensing and Bistatic Radar

### v2

@@ IV-C Clutter Removal via LoSRef-based CIR Alignment
To isolate the time-varying component, clutter is removed using a simple temporal averaging approach. Specifically, the clutter is estimated by taking the temporal average of the CIR over the observation interval, i.e.,
@@
In radar signal processing, clutter removal is a standard operation in which static components, including the direct path, are suppressed. In contrast, the proposed LoSRef-based framework differs in that the LoS-path component is first exploited as a physical delay and phase reference to achieve CIR alignment, and subsequently suppressed through clutter removal.
@@ V Experimental Analysis of OTA Phase-Referenced Channel Impulse Responses
To validate the delay-domain sensing framework described in Section IV, we conducted two representative experimental scenarios designed to capture both (i) coarse delay-domain magnitude variations induced by translational human gait and (ii) sub-wavelength, phase-sensitive variations induced by respiration.
@@ V-A Experimental Equipment
As the Tx and Rx, we used two ASUS NUC 13 Rugged-Tall devices, shown in Fig. 4, each equipped with an Intel AX210 Wi-Fi card and running PicoScenes [29] for CSI acquisition. Only a single external antenna was used at each device, resulting in a single-input single-output channel. In the following, the Tx and Rx positions are defined by the locations of the corresponding antennas. The devices were connected to a control PC via Ethernet and remotely operated to coordinate packet transmission and reception.
@@
The Tx transmitted packets using the parameters summarized in Table II. The Rx received these packets and acquired the corresponding CSI, along with the associated RSSI. The nominal packet transmission interval was set to $1\text{\,}\mathrm{ms}$; however, the actual inter-packet intervals observed at the Rx are not perfectly uniform. The characteristics of this timing irregularity are discussed in detail in Section VI.
For ground-truth position reference in the human gait experiment, HTC VIVE Tracker 3.0 [30] based on infrared optical tracking, together with SteamVR Base Station 2.0, were attached to the top of the moving human target’s head and to the Tx and Rx.
@@ V-B Frequency-Domain Preprocessing
In addition to the standard preprocessing provided by the PicoScenes Python Toolbox [29], such as interpolation of unobserved pilot subcarriers, further frequency-domain preprocessing steps were applied, partially following the procedure described in Section VI-C of [17], with additional preprocessing steps introduced in this study. After applying these preprocessing steps, the resulting frequency-domain channel representation is denoted as $H_{\mathrm{raw}}(f,t)$. The details of the preprocessing procedure are as follows.
As an outlier removal procedure, CFR samples that exhibit abrupt and extremely low RSSI are treated as outliers and excluded from further analysis. In the considered bistatic configuration, target motion affects the CFR, but is not expected to induce large changes in the overall received power, which is dominated by the LoS path. Accordingly, sudden power drops exceeding $10\text{\,}\mathrm{dB}$ are regarded as non-physical outliers since the received RSSI, reported with $1\text{\,}\mathrm{dB}$ granularity, typically varies only on the order of a few decibels.
@@
where $k\in\{$-1012$,$-1011$,\ldots,$1012$\$\} denotes the subcarrier index, with the corresponding frequency given by $f=k\varDelta f$ and the subcarrier spacing $\varDelta f=$78.125\text{\,}\mathrm{kHz}$\$.
To resolve this issue, the phase interpolation estimates the phase trend from neighboring subcarriers using the unwrapped phase. The DC subcarrier phases are then reconstructed accordingly, as confirmed by the resulting phase differences in the DC region being consistent with those of the surrounding subcarriers.
@@
In addition, as observed in Fig. 5LABEL:sub@fig:cfr_magnitude, the CFR magnitude exhibits noticeable attenuation for frequencies with $|f|>$53\text{\,}\mathrm{MHz}$\$. To quantitatively characterize this attenuation, the Tx and Rx were directly connected using coaxial cables and an attenuator, thereby forming a non-frequency-selective channel. Fig. 6 shows an example of the measured CFR, where the magnitude is normalized by the RMS value over subcarriers with $|k|\leq$680$\$. The measured CFR magnitude is approximately unity for subcarriers with $|k|\leq$680$\$, while it decreases to about $0.6$ for $|k|\geq$704$\$, with a gradual and nearly linear transition observed between these regions. To quantitatively capture this attenuation behavior, we introduce a simple piecewise attenuation model $A(k)$ in this study as
@@
which approximates the observed magnitude trend and is overlaid in Fig. 6.
Based on this model, the non-uniform frequency response is compensated by scaling the magnitude of each subcarrier by the inverse of $A(k)$. This operation effectively equalizes the spectral magnitude, resulting in a flattened and approximately uniform CFR across all subcarriers, as shown in Fig. 6.
@@ V-D1 Experimental Setup and Range–Time Representation
The Tx and Rx were placed $1.0\text{\,}\mathrm{m}$ apart in a LoS configuration, corresponding to a reference delay of $\tau_{\mathrm{ref}}=$1.0\text{\,}\mathrm{m}$/c=$3.3\text{\,}\mathrm{ns}$\$, with both antennas positioned at a height of approximately $0.8\text{\,}\mathrm{m}$ above the floor. The experimental setup and the straight-line walking trajectory representing back-and-forth translational gait motion of the human target were illustrated in Fig. 8. This simple gait scenario, involving back-and-forth translational motion along a straight-line trajectory, allows the time-varying CIR and its Doppler characteristics to be examined in a clear and interpretable manner, which will be discussed later in Section VI.
@@
Fig. 9 shows a snapshot of the calibrated CIR $h(\tau,t)$ at a representative time instant $t=$3.2\text{\,}\mathrm{s}$\$, whereas Fig. 10LABEL:sub@fig:h1_tau-t presents the corresponding CIR evolution over the entire observation interval as a delay–time heatmap. For visualization purposes, the delay axis $\tau$ is converted to propagation distance $c\tau$, yielding a range–time map. In the bistatic configuration considered here, this propagation distance represents the sum of the Tx–target and target–Rx path lengths. Following bistatic radar terminology [32], this quantity is hereafter referred to as the bistatic range $d_{\mathrm{Tx\text{--}target\text{--}Rx}}$.
Fig. 10LABEL:sub@fig:h1_tau-t shows that the strongest signal was consistently aligned at $\tau_{\mathrm{ref}}=$3.3\text{\,}\mathrm{ns}$\$ over time as a result of the LoSRef-enabled delay calibration. Moreover, the time-varying components induced by human motion are not clearly distinguishable in the raw CIR representation.
Accordingly, clutter removal described in Section IV-C is applied to obtain the residual CIR. Fig. 10LABEL:sub@fig:h_tau-t shows the resulting delay–time response, with the geometry-based bistatic range $d_{\mathrm{Tx\text{--}target\text{--}Rx}}$ overlaid for reference. A good agreement was observed between the two.
In addition, a snapshot of the residual CIR and the estimated clutter $\overline{h}(\tau)$ are shown in Fig. 9, together with the previously presented results. Although the human-induced scattering component around $\tau=$25\text{\,}\mathrm{ns}$\$ is about $20\text{\,}\mathrm{dB}$ weaker than the dominant clutter component $\overline{h}(\tau)$, the residual CIR $\tilde{h}(\tau,t)$ clearly captured the corresponding temporal variations through clutter removal.
@@ V-D2 Short-Time Phase Evolution and Doppler Interpretation
To further investigate fine-grained dynamics that are not visible at the longer time scales considered above, we focus on the short-time phase evolution of the residual CIR. Fig. 11 shows a zoomed-in view of the phase $\arg\tilde{h}($20\text{\,}\mathrm{ns}$,t)$, restricted to a duration of $0.1\text{\,}\mathrm{s}$.
As observed in Fig. 11, the phase advanced by approximately $2\pi$ every $T=$33\text{\,}\mathrm{ms}$\$, exhibiting a periodic phase evolution. This indicated that, over this short interval, the residual CIR at $\tau=$20\text{\,}\mathrm{ns}$\$ was dominated by a single complex sinusoidal component with period $T$. Accordingly, $\tilde{h}($20\text{\,}\mathrm{ns}$,t)$ can be locally approximated as $\tilde{h}($20\text{\,}\mathrm{ns}$,t)\propto\mathrm{e}^{\mathrm{j}2\pi\nu t}$, where the frequency $\nu\coloneqq 1/T=$30\text{\,}\mathrm{Hz}$\$ represents the Doppler-induced frequency offset.
@@ V-D3 Physical Interpretation in Terms of Bistatic Range Rate
This periodic phase evolution admits a direct physical interpretation in terms of bistatic range variation. A phase advance of $2\pi$ corresponds to a reduction of the bistatic range by one wavelength $\lambda$ over the same interval $T$. Defining the bistatic range rate as the time derivative of the bistatic range, $\dot{d}_{\mathrm{Tx\text{--}target\text{--}Rx}},$ the observed phase evolution implies $\dot{d}_{\mathrm{Tx\text{--}target\text{--}Rx}}=-\lambda/T.$ This relationship is consistent with the well-known bistatic Doppler relationship [32]
@@
In the present example, the positive phase slope indicates a decreasing bistatic range, corresponding to the target approaching the Tx–Rx pair. Conversely, when the target recedes, the phase slope—and hence the Doppler frequency $\nu$—becomes negative. Therefore, the sign of $\nu$ directly reveals the direction of motion.
@@
Such fine-grained, phase-coherent delay-domain observations have traditionally required dedicated channel sounders or radar systems. In contrast, the present results demonstrate that this regime can be reached using low-cost COTS Wi-Fi devices. This shift is enabled by the recent availability of wideband Wi-Fi signals with bandwidths of up to $160\text{\,}\mathrm{MHz}$, together with careful preprocessing that allows sufficiently accurate CIR estimation.
@@ V-D4 Comparison with Magnitude-Based Wi-Fi Sensing
To clarify the advantage of this phase-coherent interpretation, we now contrast it with conventional magnitude-based Wi-Fi sensing. In such approaches, temporal analysis is performed on the magnitude of the CFR, $\lvert H_{\mathrm{raw}}(f,t)\rvert$. Since the CFR consists of multiple subcarriers, dimensionality reduction is typically applied prior to temporal analysis, for example by selecting representative subcarriers (e.g., [23]) or by applying PCA (e.g., [24]).
As a representative example, Fig. 12 shows $\lvert H_{\mathrm{raw}}(f,t)\rvert$ at a representative subcarrier ($k=100$) over the same interval as Fig. 11. A periodic fluctuation with a period of approximately 33 $3\text{\,}\mathrm{ms}$ was also observed.
Unlike the phase evolution, however, these magnitude fluctuations arise from time-varying interference between human-induced scattering and static multipath components and do not preserve carrier-phase information. Consequently, while the fluctuation rate reflects motion speed, the Doppler sign—and hence whether the target is approaching or receding—cannot be inferred.
@@ V-E1 Experimental Setup and Range–Time Representation
Fig. 13LABEL:sub@fig:phase_tau_t presents a delay–time representation of the residual CIR $\tilde{h}(\tau,t)$. In this visualization, the phase of the residual CIR is shown by color, while the magnitude is represented by brightness on a logarithmic scale. A clear periodic pattern was observed around the geometry-based bistatic delay of $20\text{\,}\mathrm{ns}$, suggesting respiration-induced motion of the human target. Note that this response exhibited a noticeable spread in the delay domain. This spread arises from the finite delay resolution, $1/B=$6.3\text{\,}\mathrm{ns}$\$.
@@ V-E2 Phase Evolution and Bistatic Range Interpretation
To interpret this phase variation to physical motion, let $\varDelta d_{\text{Tx--target--Rx}}$ denote the respiration-induced variation in the bistatic range $d_{\text{Tx--target--Rx}}$. From the relationship $(2\pi/\lambda)\,\varDelta d_{\text{Tx--target--Rx}}=3\pi/2$, the corresponding path-length variation is estimated as $\varDelta d_{\text{Tx--target--Rx}}=3\lambda/4=$40\text{\,}\mathrm{mm}$\$. This value was physically reasonable for respiration-induced body-surface motion. Note that this estimated displacement does not necessarily correspond to a specific anatomical point, but rather reflects respiration-induced variations of dominant reflecting surfaces on the human body.
@@ V-E3 Comparison with Magnitude-Based Wi-Fi Sensing
It is also worth highlighting the difference between the proposed approach and commonly adopted Wi-Fi sensing methods. In many existing studies [23, 33, 34], sensing is performed on the magnitude of CFR, $\lvert H_{\mathrm{raw}}(f,t)\rvert$. Since the CFR spans multiple subcarriers, dimensionality reduction—such as subcarrier selection or PCA—is first applied, and the respiration rate is then estimated by performing an FFT along the time axis of the resulting magnitude time series.
@@
As representative examples, the time series of $\lvert H_{\mathrm{raw}}(f,t)\rvert$ at a representative subcarrier ($k=100$) and the PC2 score extracted from $\lvert H_{\mathrm{raw}}(f,t)\rvert$ are shown in Fig. 14. A clear periodic fluctuation was observed, indicating that respiration-induced motion is captured. In this sense, applying an FFT to these signals is sufficient for estimating the respiration rate, and such processing has been widely adopted in existing Wi-Fi sensing studies.
However, compared with Fig. 13LABEL:sub@fig:phase_t, these results exhibited a qualitatively different waveform with substantially larger noise-like fluctuations in magnitude. This is because these results are based on the magnitude of the CSI, which inherently mixes respiration-induced motion with other sources of magnitude variation. Moreover, magnitude-based processing does not preserve the phase corresponding to the bistatic range itself and therefore lacks a direct physical interpretation. As a result, the component associated with respiration is often identified heuristically and varies significantly across devices and environments. In particular, PC1 frequently reflects dominant static or hardware-related variations rather than motion-induced dynamics, as reported in prior studies [24, 23].
In contrast, the LoSRef-based method exploits the temporal evolution of the phase, enabling a direct physical interpretation in terms of bistatic range variations. Unlike magnitude-based sensing, this phase-based approach preserves propagation-induced information and is less sensitive to AGC-induced gain fluctuations.

### v3

@@ IV-C Clutter Removal via LoSRef-based CIR Alignment
To isolate the time-varying component, clutter is removed using a temporal averaging approach. Specifically, the clutter is estimated by taking the temporal average of the CIR over the observation interval, i.e.,
@@
In radar signal processing, clutter removal is a standard operation in which static components, including the direct path, are suppressed. In contrast, the proposed LoSRef-based framework differs in that the LoS-path component is first exploited as a delay and phase reference to achieve CIR alignment, and subsequently suppressed through clutter removal.
@@ V Experimental Analysis of OTA Phase-Referenced Channel Impulse Responses
To validate the proposed framework described in Section IV, we conducted two representative experimental scenarios designed to capture both (i) coarse delay-domain magnitude variations induced by translational human gait and (ii) sub-wavelength, phase-sensitive variations induced by respiration.
@@ V-A Experimental Equipment
As the Tx and Rx, we used two ASUS NUC 13 Rugged-Tall devices, shown in Fig. 4, each equipped with an Intel AX210 Wi-Fi card and running PicoScenes [29] for CSI acquisition. Only a single external antenna was used at each device, resulting in a single-input single-output channel. In the following, the Tx and Rx positions are defined by the locations of the corresponding antennas. The devices were connected to a control PC via Ethernet and remotely operated to coordinate packet transmission and reception. The Tx transmitted packets using the parameters summarized in Table II. The Rx received these packets and acquired the corresponding CSI, along with the RSSI. The nominal packet transmission interval was set to $1\text{\,}\mathrm{ms}$; however, the actual inter-packet intervals observed at the Rx are not perfectly uniform. The characteristics of this timing irregularity are discussed in Section VI.
For ground-truth position reference in the human gait experiment, HTC VIVE Tracker 3.0 [30] based on infrared optical tracking, together with SteamVR Base Station 2.0, were attached to the moving human target’s head and to the Tx and Rx.
@@ V-B Frequency-Domain Preprocessing
In addition to the preprocessing provided by the PicoScenes Python Toolbox [29], such as interpolation of unobserved pilot subcarriers, further frequency-domain preprocessing steps were applied, partially following the procedure described in Section VI-C of [17], with additional preprocessing steps introduced in this study. After applying these preprocessing steps, the resulting frequency-domain channel representation is denoted as $H_{\mathrm{raw}}(f,t)$. The details of the preprocessing procedure are as follows.
As an outlier removal procedure, CFR samples that exhibit abrupt and extremely low RSSI are treated as outliers and excluded from further analysis. In the considered bistatic configuration, target motion affects the CFR, but is not expected to induce large changes in the overall received power, which is dominated by the LoS path. Accordingly, sudden power drops exceeding $10\text{\,}\mathrm{dB}$ are regarded as outliers since the received RSSI, reported with $1\text{\,}\mathrm{dB}$ granularity, typically varies only on the order of a few decibels.
@@
where $k\in\{$-1012$,$-1011$,\ldots,$1012$\$\} denotes the subcarrier index, with the corresponding frequency given by $f=k\varDelta f$ and the subcarrier spacing $\varDelta f=$78.125\text{\,}\mathrm{kHz}$\$. To resolve this issue, the phase interpolation estimates the phase trend from neighboring subcarriers using the unwrapped phase. The DC subcarrier phases are then reconstructed accordingly, as confirmed by the resulting phase differences in the DC region being consistent with those of the surrounding subcarriers.
@@
In addition, as observed in Fig. 5LABEL:sub@fig:cfr_magnitude, the CFR magnitude exhibits noticeable attenuation for frequencies with $|f|>$53\text{\,}\mathrm{MHz}$\$. To quantitatively characterize this attenuation, the Tx and Rx were directly connected using coaxial cables and an attenuator, thereby forming a non-frequency-selective channel. Fig. 6 shows an example of the measured CFR, where the magnitude is normalized by the RMS value over subcarriers with $|k|\leq$680$\$. The measured CFR magnitude is approximately unity for subcarriers with $|k|\leq$680$\$, while it decreases to about $0.6$ for $|k|\geq$704$\$, with a nearly linear transition observed between these regions. To capture this attenuation behavior, we introduce a piecewise attenuation model $A(k)$ in this study as
@@
which approximates the observed magnitude trend and is overlaid in Fig. 6. Based on this model, the non-uniform frequency response is compensated by scaling the magnitude of each subcarrier by the inverse of $A(k)$. This operation effectively equalizes the spectral magnitude, resulting in an approximately uniform CFR across all subcarriers, as shown in Fig. 6.
@@ V-D1 Experimental Setup and Range–Time Representation
The Tx and Rx were placed $1.0\text{\,}\mathrm{m}$ apart in a LoS configuration, corresponding to a reference delay of $\tau_{\mathrm{ref}}=$1.0\text{\,}\mathrm{m}$/c=$3.3\text{\,}\mathrm{ns}$\$, with both antennas positioned at a height of $0.8\text{\,}\mathrm{m}$ above the floor. The experimental setup and the straight-line walking trajectory representing back-and-forth translational gait motion of the human target were illustrated in Fig. 8. This simple gait scenario, involving back-and-forth translational motion along a straight-line trajectory, allows the time-varying CIR and its Doppler characteristics to be examined in an interpretable manner, which will be discussed later in Section VI.
@@
Fig. 9 shows a snapshot of the calibrated CIR $h(\tau,t)$ at a representative time instant $t=$3.2\text{\,}\mathrm{s}$\$, whereas Fig. 10LABEL:sub@fig:h1_tau-t presents the corresponding CIR evolution over the entire observation interval as a delay–time heatmap. The delay axis $\tau$ is converted to propagation distance $c\tau$, thereby yielding a range–time map. In this bistatic configuration, it represents the bistatic range $d_{\mathrm{Tx\text{--}target\text{--}Rx}}$, i.e., the sum of the Tx–target and target–Rx path lengths [32].
Fig. 10LABEL:sub@fig:h1_tau-t shows that the strongest signal was consistently aligned at $\tau_{\mathrm{ref}}=$3.3\text{\,}\mathrm{ns}$\$ over time as a result of the LoSRef-enabled delay calibration. Meanwhile, the time-varying components induced by human motion are not clearly distinguishable in the raw CIR representation. Accordingly, clutter removal described in Section IV-C is applied to obtain the residual CIR, whose delay–time response is shown in Fig. 10LABEL:sub@fig:h_tau-t, with the geometry-based bistatic range $d_{\mathrm{Tx\text{--}target\text{--}Rx}}$ overlaid for reference. A good agreement was observed between the two. In addition, a snapshot of the residual CIR and the estimated clutter $\overline{h}(\tau)$ are shown in Fig. 9, together with the previously presented results. Although the human-induced scattering component around $\tau=$25\text{\,}\mathrm{ns}$\$ is about $20\text{\,}\mathrm{dB}$ weaker than the dominant clutter component $\overline{h}(\tau)$, the residual CIR $\tilde{h}(\tau,t)$ clearly captured the corresponding temporal variations through clutter removal.
@@ V-D2 Short-Time Phase Evolution and Doppler Interpretation
To further investigate fine-grained dynamics that are not visible at the longer time scales above, we focus on the short-time phase evolution of the residual CIR. Fig. 11 shows a zoomed-in view of the phase $\arg\tilde{h}($20\text{\,}\mathrm{ns}$,t)$.
As observed in Fig. 11, the phase advanced by approximately $2\pi$ every $T=$33\text{\,}\mathrm{ms}$\$, exhibiting a periodic phase evolution. This indicated that, the residual CIR at $\tau=$20\text{\,}\mathrm{ns}$\$ was dominated by a single complex sinusoidal component with period $T$. Accordingly, $\tilde{h}($20\text{\,}\mathrm{ns}$,t)$ can be locally approximated as $\tilde{h}($20\text{\,}\mathrm{ns}$,t)\propto\mathrm{e}^{\mathrm{j}2\pi\nu t}$, where the frequency $\nu\coloneqq 1/T=$30\text{\,}\mathrm{Hz}$\$ represents the Doppler-induced frequency offset.
@@ V-D3 Physical Interpretation in Terms of Bistatic Range Rate
This periodic phase evolution admits an interpretation in terms of bistatic range variation. A phase advance of $2\pi$ corresponds to a reduction of the bistatic range by one wavelength $\lambda$ over the same interval $T$. Defining the bistatic range rate as the time derivative of the bistatic range, $\dot{d}_{\mathrm{Tx\text{--}target\text{--}Rx}},$ the observed phase evolution implies $\dot{d}_{\mathrm{Tx\text{--}target\text{--}Rx}}=-\lambda/T.$ This relationship is consistent with the bistatic Doppler relation [32]
@@
In the example, the positive phase slope indicates a decreasing bistatic range, corresponding to the target approaching the Tx–Rx pair. Conversely, when the target recedes, the phase slope—the Doppler frequency $\nu$—becomes negative. Therefore, the sign of $\nu$ reveals the direction of motion.
@@
Such fine-grained, phase-coherent delay-domain observations have traditionally required dedicated channel sounders or radar systems. In contrast, the present results demonstrate that this regime can be reached using low-cost COTS Wi-Fi devices. This shift is enabled by the recent availability of wideband Wi-Fi signals with bandwidths of up to $160\text{\,}\mathrm{MHz}$, together with preprocessing that allows sufficiently accurate CIR estimation.
@@ V-D4 Comparison with Magnitude-Based Wi-Fi Sensing
To clarify the advantage of this approch, we now contrast it with conventional magnitude-based sensing. In these approaches, temporal analysis is performed on the magnitude of the CFR, $\lvert H_{\mathrm{raw}}(f,t)\rvert$. To reduce dimensionality, representative subcarriers are selected (e.g., [23]) PCA is applied (e.g., [24]). Fig. 12 shows $\lvert H_{\mathrm{raw}}(f,t)\rvert$ at a subcarrier ($k=100$) over the same interval as Fig. 11. A periodic fluctuation with a period of approximately 33 $3\text{\,}\mathrm{ms}$ was also observed.
Unlike the phase evolution, however, these magnitude fluctuations arise from time-varying interference between human-induced scattering and static multipath components and do not preserve carrier-phase information. Consequently, while the fluctuation rate reflects motion speed, the Doppler sign—whether the target is approaching or receding—cannot be inferred.
@@ V-E1 Experimental Setup and Range–Time Representation
Fig. 13LABEL:sub@fig:phase_tau_t presents a delay–time representation of $\tilde{h}(\tau,t)$. In this visualization, the phase of the residual CIR is shown by color, while the magnitude is represented by brightness on a logarithmic scale. A clear periodic pattern was observed around the geometry-based bistatic delay of $20\text{\,}\mathrm{ns}$, suggesting respiration-induced motion of the human target. Note that this response exhibited a spread in the delay domain. This spread arises from the finite delay resolution, $1/B=$6.3\text{\,}\mathrm{ns}$\$.
@@ V-E2 Phase Evolution and Bistatic Range Interpretation
To interpret this phase variation in terms of physical motion, let $\varDelta d_{\text{Tx--target--Rx}}$ denote the respiration-induced variation in the bistatic range $d_{\text{Tx--target--Rx}}$. From the relationship $(2\pi/\lambda)\,\varDelta d_{\text{Tx--target--Rx}}=3\pi/2$, the corresponding path-length variation is estimated as $\varDelta d_{\text{Tx--target--Rx}}=3\lambda/4=$40\text{\,}\mathrm{mm}$\$. This value was reasonable for respiration-induced body-surface motion. Note that this estimated displacement does not necessarily correspond to a specific anatomical point, but rather reflects respiration-induced variations of dominant reflecting surfaces on the human body.
@@ V-E3 Comparison with Magnitude-Based Wi-Fi Sensing
It is also worth highlighting the difference between the proposed approach and commonly adopted Wi-Fi sensing methods. In many studies [23, 33, 34], sensing is performed on the magnitude of CFR, $\lvert H_{\mathrm{raw}}(f,t)\rvert$. Since the CFR spans multiple subcarriers, dimensionality reduction—such as subcarrier selection or PCA—is first applied, and the respiration rate is then estimated by performing an FFT along the time axis of the resulting magnitude time series.
@@
As representative examples, the time series of $\lvert H_{\mathrm{raw}}(f,t)\rvert$ at a representative subcarrier ($k=100$) and the PC2 score extracted from $\lvert H_{\mathrm{raw}}(f,t)\rvert$ are shown in Fig. 14. A clear periodic fluctuation was observed, indicating that respiration-induced motion is captured. In this sense, applying an FFT to these signals is sufficient for estimating the respiration rate, and such processing has been adopted in existing Wi-Fi sensing studies.
However, compared with Fig. 13LABEL:sub@fig:phase_t, these results exhibited a qualitatively different waveform with substantially larger noise-like fluctuations in magnitude. This is because these results are based on the magnitude of the CSI, which inherently mixes respiration-induced motion with other sources of magnitude variation. Moreover, magnitude-based processing does not preserve the phase corresponding to the bistatic range itself and therefore lacks a physical interpretation. As a result, the component associated with respiration is often identified heuristically and varies across devices and environments. In particular, PC1 frequently reflects dominant static or hardware-related variations rather than motion-induced dynamics, as reported in prior studies [24, 23].
In contrast, the LoSRef-based method exploits the temporal evolution of the phase, which directly reflects bistatic range variations. Unlike magnitude-based sensing, this phase-based approach preserves propagation-induced information and is less sensitive to hardware gain fluctuations.